OpenStudy (anonymous):

rearrange the equation to make T the subject (show your working out) 5(2+t)=w

OpenStudy (anonymous):

10+5t=w => t=(w-10)/5

OpenStudy (anonymous):

10+5t=w 5t=w-10 t=(w/5)-2

OpenStudy (anonymous):

how so?:s

OpenStudy (anonymous):

(w/5)/2?

OpenStudy (anonymous):

t=(w/5)-2

OpenStudy (anonymous):

5t=w-10 and t=(w-10)/5

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Both of your answers are the same.